OpenStudy (anonymous):

write the equation of a line, with the given properties, in slope intercept form through (-6,-4) and (-4,6)

OpenStudy (anonymous):

y1-y = m (x1-x) 6--4 = 3.5 (-4--6)

OpenStudy (anonymous):

so what does y= ?

OpenStudy (anonymous):

m = slope, which i got by using the following formula: (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so y= -1x+2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

wheni plugged in the numbers that what i got

OpenStudy (anonymous):

Hold up. Figuring it out

OpenStudy (anonymous):